A car travelling at 36km/h accelerates uniformly at 2m/s^2. Find its velocity in km/h in 5s time.
1
Expert's answer
2020-06-01T14:24:58-0400

v = 36 + (2 × 5) × 3.6 = 72 (km/h)

Answer: v = 72 km/h.